OpenStudy (anonymous):

Solve. 1/x - 1/x+4 = 1/3

OpenStudy (anonymous):

are you still typing?

OpenStudy (anonymous):

Is the 1/x+4 over both the x and 4 or just the x?

OpenStudy (anonymous):

both its 1 over x+4

OpenStudy (anonymous):

\[1/x-1/(x+4) = 1/3\] Multiply both sides by x: \[1-x/(x+4) = x/3\] Write the left hand side as a single fraction: \[4/(x+4) = x/3\] Subtract x/3 from both sides: \[4/(x+4)-x/3 = 0\] Write the left hand side as a single fraction: \[(-x^2-4 x+12)/(3 (x+4)) = 0\] Multiply both sides by 3(x+4): \[-x^2-4 x+12 = 0\] Solve the quadratic equation by completing the square: Divide both sides by -1: \[x^2+4 x-12 = 0\] Add 12 to both sides: \[x^2+4 x = 12\] Add 4 to both sides: \[x^2+4 x+4 = 16\] Factor the left hand side: \[(x+2)^2 = 16\] Take the square root of both sides: \[|x+2| = 4\] Eliminate the absolute value: \[x+2 = -4 or x+2 = 4\] Subtract 2 from both sides: \[x = -6 or x+2 = 4\] Subtract 2 from both sides: \[x = -6 or x = 2\]